# The Complexity of Defining and Separating Fixpoint Formulae in Modal Logic

### v1

@@ 8. Final Discussion
For “only if” we will use the fact that $\mathsf{ML}$ enjoys the CIP over each of the considered classes $\mathbf{C}$: over all models and over $\mathbb{T}^{1}$ [Gab72], and over $\mathbb{T}^{2}$ by Proposition 12. Suppose there is an $\mathsf{ML}$-separator $\psi$ of $\varphi,\varphi^{\prime}$ and let $n$ be its modal depth. Let $\theta,\theta^{\prime}$ be respective $\mathsf{ML}^{n}$-uniform consequences of $\varphi,\varphi^{\prime}$, which exist by Proposition 7 or Proposition 10 depending on the class $\mathbf{C}$. Note that in both cases we have $\text{sig}(\theta)\subseteq\text{sig}(\varphi)$ and $\text{sig}(\theta^{\prime})\subseteq\text{sig}(\varphi^{\prime})$. Moreover, since $\psi\in\mathsf{ML}^{n}$, we have $\varphi\models\theta\models\psi\models\neg\theta^{\prime}\models\neg\varphi^{\prime}$. Thus, a Craig interpolant for $\theta,\neg\theta^{\prime}$ is a Craig separator for $\varphi,\varphi^{\prime}$.
@@
We mention some interesting open problems. First, the relative succinctness of $\mu\mathsf{ML}$ over $\mathsf{ML}$ is to the best of our knowledge open in the setting with only one accessibility relation. Second, as we have mentioned in Section 3, the separators we compute are not necessarily the logically strongest ones. The logically strongest separators of $\varphi,\varphi^{\prime}$ are precisely the $\mathsf{ML}$-uniform consequences of $\varphi$ (if they exist) and are a natural object of study. Clearly, modal definability of $\varphi$ is a sufficient condition, but not a necessary one. Let $\varphi=\psi\wedge\neg\theta_{\infty}$ and $\varphi^{\prime}=\neg\psi$ for some $\psi\in\mathsf{ML}$ and $\theta_{\infty}$ from Example 3. Then $\varphi$ is not equivalent to a modal formula, but $\psi$ is a strongest separator. It would also be interesting to investigate separability of $\mu\mathsf{ML}$-formulae by other fragments of $\mu\mathsf{ML}$, such as safety formulae, for which definability has been studied in [LQ15] and to extend the $\mu$-calculus by other operators, e.g., nominals or inverse modalities.
Finally, it would be interesting to extend our findings in the case study on the graded $\mu\mathsf{ML}$ in Section 7. As already pointed out, there are two natural directions for this: first one could study the complexity under binary coding of grades, and second one could tackle the separator construction problem. We conjecture that the first problem can be solved by combining our techniques with the automata constructions from [KSV02], which lead to optimal complexity for satisfiability of $\mathsf{grML}$ under binary coding of grades. The computation problem might be much more challenging. Indeed, it was observed recently that the construction of $\mathsf{ML}$ separators of $\mathsf{grML}$-formulae already requires significantly new ideas [JKW25].

### v2

@@ 8. Final Discussion
For “only if” we will use the fact that $\mathsf{ML}$ enjoys the CIP over each of the considered classes $\mathbf{C}$: over all models and over $\mathbb{T}^{1}$ [18], and over $\mathbb{T}^{2}$ by Proposition 12. Suppose there is an $\mathsf{ML}$-separator $\psi$ of $\varphi,\varphi^{\prime}$ and let $n$ be its modal depth. Let $\theta,\theta^{\prime}$ be respective $\mathsf{ML}^{n}$-uniform consequences of $\varphi,\varphi^{\prime}$, which exist by Proposition 7 or Proposition 10 depending on the class $\mathbf{C}$. Note that in both cases we have $\text{sig}(\theta)\subseteq\text{sig}(\varphi)$ and $\text{sig}(\theta^{\prime})\subseteq\text{sig}(\varphi^{\prime})$. Moreover, since $\psi\in\mathsf{ML}^{n}$, we have $\varphi\models\theta\models\psi\models\neg\theta^{\prime}\models\neg\varphi^{\prime}$. Thus, a Craig interpolant for $\theta,\neg\theta^{\prime}$ is a Craig separator for $\varphi,\varphi^{\prime}$.
@@
We mention some interesting open problems. First, the relative succinctness of $\mu\mathsf{ML}$ over $\mathsf{ML}$ is to the best of our knowledge open in the setting with only one accessibility relation. Second, as we have mentioned in Section 3, the separators we compute are not necessarily the logically strongest ones. The logically strongest separators of $\varphi,\varphi^{\prime}$ are precisely the $\mathsf{ML}$-uniform consequences of $\varphi$ (if they exist) and are a natural object of study. Clearly, modal definability of $\varphi$ is a sufficient condition, but not a necessary one. Let $\varphi=\psi\wedge\neg\theta_{\infty}$ and $\varphi^{\prime}=\neg\psi$ for some $\psi\in\mathsf{ML}$ and $\theta_{\infty}$ from Example 3. Then $\varphi$ is not equivalent to a modal formula, but $\psi$ is a strongest separator. It would also be interesting to investigate separability of $\mu\mathsf{ML}$-formulae by other fragments of $\mu\mathsf{ML}$, such as safety formulae, for which definability has been studied in [28] and to extend the $\mu$-calculus by other operators, e.g., nominals or inverse modalities.
Finally, it would be interesting to extend our findings in the case study on the graded $\mu\mathsf{ML}$ in Section 7. As already pointed out, there are two natural directions for this: first one could study the complexity under binary coding of grades, and second one could tackle the separator construction problem. We conjecture that the first problem can be solved by combining our techniques with the automata constructions from [27], which lead to optimal complexity for satisfiability of $\mathsf{grML}$ under binary coding of grades. The computation problem might be much more challenging. Indeed, it was observed recently that the construction of $\mathsf{ML}$ separators of $\mathsf{grML}$-formulae already requires significantly new ideas [20].